OpenStudy (anonymous):

Use PFD to solve: \[\int\limits_{}^{}\frac{1}{(1+ x^2)(x)}dx\]

OpenStudy (anonymous):

PFD => Partial fraction decomposition

OpenStudy (anonymous):

@Mertsj @Zarkon @dpaInc @jim_thompson5910 @lgbasallote @KingGeorge @AccessDenied @SmoothMath @eliassaab @asnaseer @.Sam. @kropot72 @eigenschmeigen @satellite73 @amistre64 @TuringTest @shivam_bhalla @experimentX

OpenStudy (anonymous):

----- I do not know why it is saying I am typing a reply

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+1%2F%5B%281%2Bx%5E2%29%28x%29%5D I got confused at wolf's substitution

OpenStudy (anonymous):

pretty extensive list there.... :)

OpenStudy (anonymous):

\[\frac{1}{x \left( 1 + x^2 \right)} = \frac{A}{x} + \frac{Bx + C}{1 + x^2}\]\[1 = (A + B)x^2 + Cx + A\]\[A + B = 0\]\[C = 0\]\[A = 1\]solve for B\[B = -1\]and then the PFD is finished\[\frac{1}{x\left(1 + x^2\right)} = \frac{1}{x} - \frac{x}{1 + x^2}\]integrate it and its done

OpenStudy (anonymous):

Got it.